OpenStudy (anonymous):

please help,what is the equation of the parabola in vertex form ? x^2+6x-y+3=0

OpenStudy (anonymous):

x^2+6x-y+3=0 y = x^2+6x+3 = x^2 + 6x + 9 - 9 + 3 = (factorize the first three terms and simplify the last two terms)

OpenStudy (anonymous):

thanks but im still confused its multiple choice

OpenStudy (anonymous):

so itd be (x+3)^2=(y+6)?

OpenStudy (anonymous):

How do you factorize x^2 + 6x + 9?

OpenStudy (tkhunny):

The best approach for a multiple choice question is to work the problem in its entirety, get the right answer, and only then look at the choices. It's not the fastest way to proceed, but it is the most sure.

OpenStudy (anonymous):

thanks?

OpenStudy (anonymous):

x^2 goes into 6x twice and then 6 goes into 9 once,i suck at math sorry

OpenStudy (anonymous):

(a+b)^2 = a^2 + 2ab + b^2 Use this to factorize x^2 + 6x + 9

OpenStudy (anonymous):

Behold !